OpenStudy (anonymous):

SO I was curious, what is a time where you would add exponents and then also when you would multiply exponents..... I was kinda confused and was wanting examples of each, help please?

OpenStudy (anonymous):

add them when you are multiplying \[\huge 2^5\times 2^3=2^{5+3}=2^8\]

OpenStudy (anonymous):

\[\huge b^n\times b^m=b^{n+m}\]

OpenStudy (anonymous):

multiply when you are raising to powers \[\huge \left(2^5\right)^3=2^{5\times 3}=2^{15}\]

OpenStudy (anonymous):

\[\huge \left(b^m\right)^n=b^{m\times n}\]

OpenStudy (anonymous):

Oh, that makes a lot more sense. What I was given through school just said it, and didn't show. Thank you @satellite73